# TDE-3: An improved prior for optical flow computation in spiking neural networks

### v1

@@ Abstract
Motion detection is a primary task required for robotic systems to perceive and navigate in their environment. Proposed in the literature bioinspired neuromorphic Time-Difference Encoder (TDE-2) combines event-based sensors and processors with spiking neural networks to provide real-time and energy-efficient motion detection through extracting temporal correlations between two points in space. However, on the algorithmic level, this design leads to loss of direction-selectivity of individual TDEs in textured environments. Here we propose an augmented 3-point TDE (TDE-3) with additional inhibitory input that makes TDE-3 direction-selectivity robust in textured environments. We developed a procedure to train the new TDE-3 using backpropagation through time and surrogate gradients to linearly map input velocities into an output spike count or an Inter-Spike Interval (ISI). Our work is the first instance of training a spiking neuron to have a specific ISI. Using synthetic data we compared training and inference with spike count and ISI with respect to changes in stimuli dynamic range, spatial frequency, and level of noise. ISI turns out to be more robust towards variation in spatial frequency, whereas the spike count is a more reliable training signal in the presence of noise. We performed the first in-depth quantitative investigation of optical flow coding with TDE and compared TDE-2 vs TDE-3 in terms of energy-efficiency and coding precision. Results show that on the network level both detectors show similar precision (20°angular error, 88% correlation with ground truth). Yet, due to the more robust direction-selectivity of individual TDEs, TDE-3 –based network spike less and hence is more energy-efficient. Reported precision is on par with model-based methods but the spike-based processing of the TDEs provides allows more energy-efficient inference with neuromorphic hardware.
@@ 1 Introduction
The efficiency of event-based processing in solving optical flow is especially well utilized by spiking neural networks (SNNs) - computational abstractions of biological neurons that transmit signals in an all-or-none spiking fashion only when the internal neuronal activation reaches a certain threshold [3](Figure 1). Indeed, asynchronous spike seems like a perfect match to asynchronous all-or-none events. Studies showed that when it comes to optical flow SNNs can perform on par [4] or even better [5, 6] than state-of-the-art traditional neural networks, while requiring up to 50-times fewer parameters [5] and promising up to $10^{3}$ decrease in compute energy [7]. With the development of specially designated to processing SNNs neuromorphic processors [8], event-based optical flow estimation with SNNs reached a significant milestone of technological maturity - implementation for vision-based navigation onboard a drone [9].
@@
Although fundamentally successful estimation of optical flow by SNNs is due to them being universal function approximators [10, 11, 3], the choice of proper priors, such as neural architecture and connectivity on hardware and algorithmic levels, can boost their performance. For example in image classification tasks, the introduction of randomly initialized and non-trainable convolutional kernels improves the classification accuracy compared to fully connected networks [12] simply because convolutional kernels extract a spatial pattern. Similarly, proper priors to extract spatio-temporal patterns could enhance algorithms of the optical flow estimation from event-based data.
@@
A good choice of computational prior for the optical flow estimation from event-based data would be ubiquitously employed in the animal kingdom correlation-based motion detector [13]. As its name implies, this detector extracts local motion direction and velocity (i.e. optical flow) based on the temporal correlations between different points in space [13]. Such correlation-based estimates were shown to closely approach the optimal way to extract motion signals in conditions with low SNR [14, 15]. Event-based cameras are prone to noisy outputs [16] due to the environment (low contrast and light intensity [17]), the temperature [18, 19], the junction leakage [19, 20], the hardware limitations (manufacturing mismatch [1]), and the very differential nature of processing which amplifies high-frequent noise [20, 17, 21]. Consequently, a correlation-based motion detector is a good match for event-based data.
@@
Bio-inspired correlation-based methods for optical flow estimation with SNNs generated long-lasting interest in event-based vision and neuromorphic computing communities with over a decade of relevant research on both software [22, 23, 17, 24, 25] and hardware levels [26, 27, 28, 29, 30, 31, 32, 33]. In particular for our study, we are interested in the Time-Difference Encoder (TDE) [27], inspired by the classical model of insects motion detector [34]. The TDE detects motion by comparing signals at two points in space: the facilitator input provides a time-decaying gain, whereas the trigger event converts it to an input current impulse to a Current-Based Leaky-Integrate-and-Fire (CuBa LIF) neuron, which integrates it in its membrane potential and converts it into a firing activity as shown in Figure 1. The TDE direction-selectivity is achieved since it produces output spikes only when the trigger event arrives after the facilitator one, as otherwise there is no gain to be converted to firing activity. The TDE was successfully implemented on FPGA [31], digital neuromorphic hardware such as SpiNNaker [30], Loihi  [^1] TDE implementation on Intel Loihi: https://github.com/intel-nrc-ecosystem/models and Loihi 2 [33], and custom mixed-signal neuromorphic hardware [27]. Its potential was demonstrated in robotics applications for collision avoidance [27, 28]. The strength of the underlying computational primitive was also explored in tasks unrelated to vision, such as keyword spotting [35].
@@
Despite its appeal, the classical two-point TDE has a notable deficiency: in highly textured environments, it loses direction-selectivity because motion in a non-preferred direction provides the TDE with residual activity in the gain compartment, which can be subsequently triggered by another edge moving in the non-preferred direction as shown in Figure 1. Given that successful event-based navigation requires events that are produced by the apparent motion of textures[9], loss of direction-selectivity makes classical TDE-2 unsuitable for vision-based navigation tasks that require precision in optical flow estimation. Of course, this loss in direction-selectivity can be partially compensated by either max pooling [33], Winner-Take-All mechanism [27, 28] or subtraction of the velocities estimated by opposingly tuned detectors. Yet in that case, energy-efficiency of the circuitry would be obstructed as individual TDEs would still respond to the non-preferred motion.
@@
To solve this problem, we propose to follow recent findings in neuroscience [36] and to augment the TDE with a third, inhibitory input. The activation of this inhibitory input resets the gain to zero and eliminates the residual activity to retain the TDE direction-selectivity (Figure 1). Consequently, even in highly textured environments, individual TDEs retain direction selectivity, and the entire circuit spikes much less, becoming more energy-efficient. To distinguish the classical 2-point TDE from the augmented 3-point TDE, from now on we will call them TDE-2 and TDE-3, respectively.
@@
Upon fixing issues related to detector architecture, we perform a more in-depth study of the TDE coding of velocity with synthetic and real-world data. Specifically, there are four major contributions of this paper:
@@
    We present a novel, more robust TDE-3 and highlight its advantage over the TDE-2 on the level of direction-selectivity of a single detector (Figures 1 & 2).
@@
    We develop a pipeline for the supervised training of the TDE using Back-Propagation Through Time (BPTT) and surrogate gradients [37, 38] to linearly map input velocities into an output spike count (Figure 9) or an inter-spike interval (ISI) between the first and the second spike (Figure 10) within both wide (10-fold, Figure 3) and narrow (1.5-fold, Figure 4) dynamic range with precision of up to 2% (Table 1). To the best of our knowledge, this work represents the first attempt at training spiking neurons to have a particular ISI.
@@
    We compare the robustness of velocity coding with TDE to noise (Figure 6, Table 2) and spatial frequency (Figure 5) composition of the scene - two major challenges for velocity coding with correlation-based motion detectors. Our results show that spike count is more robust to noise, while ISI is less affected by changes in spatial frequency and provides faster velocity inference.
@@
    We performed first-ever quantitative characterization of resolution, robustness and energy efficiency of encoding real-world optical flow by TDEs (Figures 7, Figures 8 and Tables 3, 4). Indeed, although Schoepe et al. [28] quantitatively assesed dependence of the TDE output on velocity of real-world moving gratings, they did so on the network level and by measuring activity over the entire sequence rather than based on individual detector activations. Our results show that the TDE performs on par with model-based methods of local motion estimation reported in the literature ([39, 2] ( directional error < 20°), but has the advantage of being naturally compatible with spike-based neuromorphic hardware. Furthermore, the TDE-3 is more energy-efficient than the TDE-2, requiring up to 2×2$\times$2 × fewer spikes (Table 3). To the best of our knowledge, it is the first-ever quantitative assessment of optical flow estimation from real-world data with TDE.
@@
2 Theory of Time-Difference Encoders
Before discussing main findings of the paper it is very important to develop firm understanding of the TDE-2 and TDE-3 computational principle on the level of direction-selectivity and velocity coding to understand features and limitations of the systems under investigation.
@@ 2.1 Computational primitive and detector architecture
The TDE-2 computational principle is to extract temporal correlations at two spatial locations: facilitator and trigger. A spike input to the facilitator provides the detector with a time-decaying gain, while a spike input to the trigger converts the gain into an excitatory current which charges a CuBa-LIF neuron and leads to spiking [27] (Figure 1). Essentially, the time-decaying gain extracts temporal correlations as the neuron effectively reports the coincidence of inputs to the facilitator and trigger within a temporal window of the gain decay and scales the current according to the delay between activation of the facilitator and trigger. Hence, faster motion leads to a higher number of spikes with shorter intervals between them and vice versa.
The direction-selectivity of the neuron is ensured by the combination of the two factors. Its activation in the anti-preferred or null direction (ND) is prevented by the fact that when activation of the trigger precedes the activation of the facilitator, there is no gain to convert to current (Figure 1A center). In the case of the motion in the orthogonal direction (OD), the facilitator and trigger are activated simultaneously. In the hardware, detector firing is prevented by the delay between the processing of the trigger and gain. Here the delay is realized through the sequence of computational steps: TDE-2 first processes trigger and then gain. Hence, in the case of OD motion, there is also no gain to be converted to the current at the moment of trigger activation.
@@
Our TDE-3 solves the problem of the residual gain by incorporating a third, inhibitory input. Relative to the gain, the inhibitory input is located on the opposite side of the detector and upon activation resets the gain to zero (Figure 1). As a result of this arrangement, when secondary stimuli will be moving in ND, it will first appear at the inhibitory flank. This will remove any residual gain activity, such that when the stimulus activates the trigger, there again will be no gain to be converted to current. In the case of motion in OD, detector activation will be prevented by simultaneous processing of facilitator and inhibitory inputs (assuming stimuli wide enough to cover 3 inputs) such that the OD motion will leave no trace in gain activity. Thus, the addition of the inhibitory flank allows the detector to retain direction-selectivity in textured environments (Figure 2). As a result of such more robust direction-selectivity of individual TDEs, the entire network can spike up to two times less (Table 3).
@@ 2.2 Velocity coding with TDE
When the moving stimulus passes the TDE, two properties of the emitted spike train allow downstream circuitry to estimate velocity: the number of spikes and ISI [27, 31]. The spike count is directly proportional to the input velocity, while the ISI is inversely proportional. Indeed, fast motion means a short temporal delay between the activation of the gain and its conversion to current. Consequently, the current is large and the number of emitted spikes is high. Conversely, such vigorous firing means that spikes closely follow each other and the interval between them is small [27]. Of special interest is ISI between the first two spikes as it allows very fast velocity interest [27].
@@
From the perspective of a coding capacity of a single neuron and regardless of the inference method, all-or-none spiking is rather a liability than an asset as it introduces a latency vs. dynamic range dilemma because spiking occurs in time [40]. For example, to discriminate input velocities within a range of, e.g., 100 different values would require a coding window of at least 100 timesteps either to count spikes or to measure interspike intervals. Even with a very short timestep of 1 ms, the lag becomes substantial for edge applications like robotics, where neuromorphic hardware is needed the most and low latency is paramount. Thus, to avoid impractical latency, one can either encode narrow range of input velocities with high resolution or wide range of input velocities with low resolution (Figures 3, 4). Therefore, a strategy to reconcile latency and dynamic range requirements would be to use each detector within a relatively narrow range of velocities, while boosting the dynamic range of the entire network by using detectors with different spacing between detector inputs.
The coding of velocity by the TDE is additionally troubled by the noise and spatial frequency of a scene. Noise leads to the non-stimulus-related activation of detectors. Spatial frequency bias detector responses as edges closely following each other will lead to stronger TDE activation than a single edge. In the results section, we compare spike count and ISI-based inference in terms of their robustness in various noise (Figure 6, Table 2) and spatial frequency (Figure 5) conditions.
@@ 3 Experiments and results
The experiments that we performed can be classified into three categories: (1) investigation of the robustness of the direction-selectivity of individual TDE-3 and TDE-2 (Figure 2), (2) TDEs training to linearly map input velocities into output (Figures 3-6), and (3) application of the TDEs for optical flow inference from real-world data (Figures 7, 8).

### v2

@@ Abstract
Motion detection is a primary task required for robotic systems to perceive and navigate in their environment. Proposed in the literature bioinspired neuromorphic Time-Difference Encoder (TDE-2) combines event-based sensors and processors with spiking neural networks to provide real-time and energy-efficient motion detection through extracting temporal correlations between two points in space. However, on the algorithmic level, this design leads to a loss of direction-selectivity of individual TDEs in textured environments. Here we propose an augmented 3-point TDE (TDE-3) with additional inhibitory input that makes TDE-3 direction-selectivity robust in textured environments. We developed a procedure to train the new TDE-3 using backpropagation through time and surrogate gradients to linearly map input velocities into an output spike count or an Inter-Spike Interval (ISI). Using synthetic data we compared training and inference with spike count and ISI with respect to changes in stimuli dynamic range, spatial frequency, and level of noise. ISI turns out to be more robust towards variation in spatial frequency, whereas the spike count is a more reliable training signal in the presence of noise. We performed in-depth quantitative investigation of optical flow coding with TDE and compared TDE-2 vs TDE-3 in terms of energy-efficiency and coding precision. Results show that on the network level, both detectors show similar precision (20°angular error, 88% correlation with ground truth). Yet, due to the more robust direction-selectivity of individual TDEs, TDE-3 –based network spike less and hence is more energy-efficient. Reported precision is on par with model-based methods but the spike-based processing of the TDEs provides allows more energy-efficient inference with neuromorphic hardware.
@@ 1 Introduction
The efficiency of event-based processing in the estimation of optical flow is especially well utilized by spiking neural networks (SNNs) - computational abstractions of biological neurons that transmit signals in an all-or-none spiking fashion only when the internal neuronal activation reaches a certain threshold [3](Figure 1). Indeed, asynchronous all-or-none spikes seem like a perfect match to asynchronous all-or-none events due to the similar nature of these signals. Studies showed that when it comes to optical flow SNNs can perform on par [4] or even better [5, 6] than state-of-the-art traditional neural networks, while requiring up to 50-times fewer parameters [5] and promising up to $10^{3}$ decrease in compute energy [7]. With the development of specially designated to processing SNNs neuromorphic processors [8], event-based optical flow estimation with SNNs reached a significant milestone of technological maturity - implementation for vision-based navigation onboard a drone in a fully neuromorphic pipeline [9].
@@
Although theoretically, SNNs are capable of performing a wide variety of tasks [10, 11, 12, 13], their actual performance strongly depends on the choice of proper priors such as neural network architecture and connectivity on the hardware and algorithmic levels. For example, in image classification tasks, the introduction of randomly initialized and non-trainable convolutional kernels improves the classification accuracy compared to fully connected networks [14] simply because convolutional kernels extract a local spatial pattern. Similarly, proper priors to extract spatio-temporal patterns could enhance algorithms of the optical flow estimation from event-based data.
@@
A good computational prior for optical flow estimation from event-based data could be the correlation-based motion detector, which is ubiquitously employed in the animal kingdom [15]. As its name suggests, this detector extracts local motion direction and velocity (i.e. optical flow) based on the temporal correlations between different points in space [15]. Such correlation-based estimates were shown to closely approach the optimal way to extract motion signals in conditions with low Signal-to-Noise Ratio (SNR)[16, 17]. Event-based cameras, meanwhile, are prone to noisy outputs [18] due to the environment (low contrast and light intensity [19]), the temperature [20, 21], the junction leakage [21, 22], the hardware limitations (manufacturing mismatch [1]), and the very differential nature of processing which amplifies high-frequency noise [22, 19, 23]. Given such low SNR output, a correlation-based motion detector is well-suited to process event-based data.
@@
Bio-inspired correlation-based methods for optical flow estimation with SNNs generated long-lasting interest in event-based vision and neuromorphic computing communities with over a decade of relevant research on both software [24, 25, 19, 26, 27] and hardware levels [28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38]. In particular, for our study, we are interested in the inspired by the classical model of the insect motion detector [39] Time-Difference Encoder (TDE) [29] due to its circuit and algorithmic simplicity [29]. The TDE detects motion by comparing signals at two points in space: input to the facilitator provides a time-decaying gain, whereas input to the trigger converts gain to an input current impulse to a Current-Based Leaky Integrate-and-Fire (CuBa LIF) neuron, which integrates it in its membrane potential and converts it into a firing activity as shown in Figure 1. The TDE tuning to the preferred direction of motion (PD) is achieved since it produces output spikes only when the trigger event arrives after the facilitator one, as otherwise there is no gain to be converted to firing activity. The TDE was successfully implemented on FPGA [33], digital neuromorphic hardware such as SpiNNaker [32], Loihi  [^1] TDE implementation on Intel Loihi: https://github.com/intel-nrc-ecosystem/models and Loihi 2 [35], and custom mixed-signal neuromorphic hardware [29]. Its potential was demonstrated in robotics applications for collision avoidance [29, 30] and estimation of ego-motion [40]. The strength of the underlying computational primitive was also explored in tasks that are not related to vision, such as spotting keywords [41].
@@
Despite its appeal, the classical two-point TDE (TDE-2) has a notable deficiency: in highly textured environments, it loses direction selectivity because motion in a non-preferred direction provides the TDE with residual activity in the gain compartment, which can be subsequently triggered by another edge moving in the non-preferred direction as shown in Figure 1. Given that successful event-based navigation requires events that are produced by the apparent motion of textures[9], loss of direction-selectivity makes classical TDE-2 unsuitable for vision-based navigation tasks that require precision in optical flow estimation. Of course, this loss in direction-selectivity can be partially compensated by either max pooling [35], Winner-Take-All mechanism [29, 30] or subtraction of the velocities estimated by opposingly tuned detectors. Yet in that case, the energy-efficiency of the circuitry would be obstructed as individual TDEs would still respond to the non-preferred motion.
@@
To solve this problem, we propose to follow recent findings in neuroscience[42] and to augment the TDE with a third, inhibitory input. While the classical model of elementary motion sensitivity in insects [39, 15] estimated temporal correlations from two spatial locations, Haag et al. [42] showed that insect elementary motion detectors (so-called T4 and T5 cells) pools signals from three spatial locations with one of them being inhibitory and activated by the motion in ND. In our implementation, activation of this inhibitory input resets the gain to zero and eliminates the residual activity to retain the TDE direction-selectivity (Figure 1). Consequently, even in highly textured environments, individual TDEs retain direction selectivity, and the entire circuit spikes much less, becoming more energy-efficient. We propose to call this detector TDE-3.
@@
Upon fixing issues related to detector architecture, we performed a more in-depth study of the TDE coding of velocity with synthetic and real-world data. Specifically, there are five major contributions of this paper:
@@
    We present a more robust TDE-3 and highlight its advantage over the TDE-2 on the level of direction-selectivity of a single detector (Figures 1 & 2).
@@
    We develop a pipeline for the supervised training of the TDE using Back-Propagation Through Time (BPTT) and surrogate gradients [43, 44] to linearly map input velocities into an output spike count (Figure 12) or an inter-spike interval (ISI) between the first and the second spike (Figure 13) within both wide (10-fold, Figure 3) and narrow (1.5-fold, Figure 4) dynamic range with precision of up to 2% (Table 1).
@@
    We develop a novel approach for training SNNs to have a specific ISI using BPTT. While most other methods [45, 46, 47] model the relationship between input magnitude with differentiable function than use BPTT, we measure ISI from the amplitude of low-pass filtered spike train to use BPTT with surrogate gradient (see Section 5.2.2 for the details). The theoretical advantage of our method is that it does not require estimation of the relationship between spike timing and input magnitude, allowing work with more conventional activation functions.
@@
    We compare the robustness of velocity coding with TDEs to noise (Figures 6, 7) and spatial frequency (Figure 5) composition of the scene - two major challenges for velocity coding with correlation-based motion detectors. Our results show that spike count is more robust to noise, while ISI is less affected by changes in spatial frequency and provides faster velocity inference.
  5. 5.
    We performed the first-ever quantitative characterization of resolution, robustness, and energy efficiency of encoding real-world optical flow by TDEs (Figures 10, Figures 11 and Tables 2, 3). Indeed, although Schoepe et al. [30] quantitatively assessed the dependence of the TDE output on the velocity of real-world moving gratings, they did so on the network level and by measuring activity over the entire sequence rather than based on individual detector activations. Our results show that the TDE performs on par with model-based methods of local motion estimation reported in the literature ([48, 2] ( directional error < 20°), but has the advantage of being naturally compatible with spike-based neuromorphic hardware. Furthermore, the TDE-3 is more energy-efficient than the TDE-2, requiring up to 24×2-4$\times$2 - 4 × fewer spikes (Table 2, Supplementary Information).
@@
2 Time-Difference Encoders: Principles of Operation
Before discussing the main findings of the paper it is important to understand the TDE-2 and TDE-3 computational principles on the level of direction-selectivity and velocity coding to get insight into the features and limitations of the systems under investigation.
@@ 2.1 Computational primitive and detector architecture
The TDE-2 computational principle is to extract temporal correlations at two spatial locations: facilitator and trigger. A spike input to the facilitator provides the detector with a time-decaying gain, while a spike input to the trigger converts the gain into an excitatory current which charges CuBa-LIF neuron and leads to spiking [29] (Figure 1). Essentially, the time-decaying gain extracts temporal correlations as the neuron reports the coincidence of inputs to the facilitator and trigger within a temporal window of the gain decay and scales the current according to the delay between activation of the facilitator and trigger. Hence, faster motion leads to a higher number of spikes with shorter intervals between them and vice versa.
The direction-selectivity of the neuron is ensured by the combination of the two factors. Its activation in the direction anti-parallel to PD, also known as null direction (ND), is prevented since when activation of the trigger precedes the activation of the facilitator, there is no gain to convert to current (Figure 1A center). In the case of the motion in the direction orthogonal to PD (OD), the facilitator and trigger are activated simultaneously. In the hardware, detector firing is prevented by the delay between the processing of the trigger and gain. Here the delay is realized through the sequence of computational steps: TDE-2 first processes trigger and then gain. Hence, in the case of OD motion, there is also no gain to be converted to the current at the moment of trigger activation.
@@
Our TDE-3 solves the problem of the residual gain by incorporating a third, inhibitory input. Relative to the gain, the inhibitory input is located on the opposite side of the detector and upon activation resets the gain to zero (Figure 1). As a result, when secondary stimuli are moving in the ND, they will first appear at the inhibitory flank. This will remove any residual gain activity, such that when the stimulus activates the trigger, there again will be no gain to be converted to current. In the case of motion in the OD, detector activation will be prevented by simultaneous processing of facilitator and inhibitory inputs (assuming stimuli wide enough to cover 3 inputs) such that the OD motion will leave no trace in gain activity. Thus, the addition of the inhibitory flank allows the detector to retain direction-selectivity in textured environments (Figure 2). As a consequence of such more robust direction-selectivity of individual TDEs, the entire network can spike up to 2-4 times less (Table 2, Supplementary Information Figure S2).
@@ 2.2 Velocity coding with TDE
When the moving stimulus passes the TDE, two properties of the emitted spike train allow downstream circuitry to estimate velocity: the number of spikes and ISI [29, 33]. The spike count is directly proportional to the input velocity, while the ISI is inversely proportional. Indeed, fast motion means a short temporal delay between the activation of the gain and its conversion to current. Consequently, the current is large and the number of emitted spikes is high. Conversely, such vigorous firing means that spikes closely follow each other and the interval between them is small [29]. Of special interest is ISI between the first two spikes as it allows very fast velocity inference [29].
@@
From the perspective of a single neuron’s coding capacity, and regardless of the inference method, all-or-none spiking is more of a liability than an asset. This is because it introduces a latency vs. dynamic range dilemma, as spiking occurs in time [49]. For example, to discriminate input velocities within a range of, e.g., 100 different values would require a coding window of at least 100 timesteps either to count spikes or to measure interspike intervals. Even with a very short timestep of 1 ms, the lag becomes substantial for edge applications like robotics, where neuromorphic hardware is needed the most and low latency is paramount. Thus, to avoid impractical latency, one can either encode a narrow range of input velocities with high resolution or a wide range of input velocities with low resolution (Figures 3, 4). Therefore, a strategy to reconcile latency and dynamic range requirements would be to use each detector within a relatively narrow range of velocities, while boosting the dynamic range of the entire network by using detectors with different spacing between detector inputs.
The coding of velocity by the TDE is additionally troubled by the noise and spatial frequency of a scene. Noise leads to the non-stimulus-related activation of detectors. Spatial frequency bias detector responses as edges closely following each other will lead to stronger TDE activation than a single edge. In the results section, we compare spike count and ISI-based inference in terms of their robustness in various noise (Figures 6, 7) and spatial frequency (Figure 5) conditions.
@@ 3 Experiments and results
The experiments that we performed can be classified into three categories: (1) investigation of the robustness of the direction-selectivity of individual TDE-3 and TDE-2 (Figure 2), (2) TDEs training to linearly map input velocities into output (Figures 3-6), and (3) application of the TDEs for optical flow inference from real-world data (Figures 10, 11).